OpenStudy (anonymous):

When 10.6g of a nonvolatile substance is dissolved in 740g of ether, it boiling point is raised 0.284. What is the molar mass of the substance?Kb=2.11C/m for ether.

OpenStudy (anonymous):

I calculated the rate constant k by dividing 0.284/2.11 and got 1.3459*10^-1m

OpenStudy (anonymous):

But i don't know what to do next.

OpenStudy (aaronq):

you're have to use the boiling point elevation equation

OpenStudy (aaronq):

you*

OpenStudy (anonymous):

yeah, and I got the value of molality, now what??

OpenStudy (anonymous):

\[I did: \Delta T _{b}=K _{b}m\] \[\frac{ 0.284 }{ 2.11 }=m\]

OpenStudy (aaronq):

how is molality defined?

OpenStudy (aaronq):

(mathematically)

OpenStudy (anonymous):

moles of solute/kg of solvent

OpenStudy (anonymous):

I would be able to go forward if I had the molar mass of something, but I dont

OpenStudy (aaronq):

the molar mass IS what you're asked to find

OpenStudy (anonymous):

yeah, i meant the molar mass of ether to convert stuff, I dont know, I feel I'm stuck

OpenStudy (aaronq):

how is moles defined? substitute that into the molality equation

OpenStudy (anonymous):

moles? oohh!! 6.022*10^23??

OpenStudy (aaronq):

no no you're thinking too deep, molality = moles of solute/ Kg of solvent substitute the definition into this equation

OpenStudy (aaronq):

definition of moles that is lol

OpenStudy (anonymous):

so I have 1.3459*10^-1 moles solute/1 kg (or 1000g) of solvent

OpenStudy (aaronq):

nope nope you're thinking about how much 1 mole is we're talking about moles=mass/molar mass

OpenStudy (aaronq):

you can't possibly know how many moles you have if you don't know the molar mass

OpenStudy (anonymous):

yeah, so how can I convert? I'm sorry! :S

OpenStudy (aaronq):

haha it's okay. okay so molality = moles of solute/ Kg of solvent moles = mass/ molar mass, right? substitute the equation for moles into the molality equation

OpenStudy (anonymous):

\[\frac{ \frac{ mass }{ molar mass } solute }{ 1kg solvent }\]

OpenStudy (anonymous):

[mass/molar mass] solute/kg solvent

OpenStudy (anonymous):

i am seriosuly trying...

OpenStudy (aaronq):

sorry i typed some stuff wrong, this is what i meant: you have 4 variables (including molality) and 3 knowns, 1 unknown molality = (mass/molarmass)/Kg of solvent

OpenStudy (anonymous):

\[1.3459*10^{-1} = \frac{ 10.6 g}{ molar mass } / 7.4*^{-3}\]

OpenStudy (anonymous):

????

OpenStudy (aaronq):

isolate molar mass

OpenStudy (aaronq):

you have the mass of solvent wrong it's 0.740 Kg

OpenStudy (anonymous):

I get 106.4g/mol

OpenStudy (aaronq):

yep thats right

OpenStudy (anonymous):

thanks!!!

OpenStudy (aaronq):

no problem !